# Efficient Rule Induction by Ignoring Pointless Rules

### v2

@@ 1 Introduction
Given these inputs, we might want to learn a rule such as:
@@
Although useful, we argue that $r_{2}$ and $r_{3}$ are pointless because odd(A) implies int(A), so we can remove int(A) whilst preserving semantics. We call such rules reducible rules.
@@
The novelty of this paper is to show that ignoring reducible and indiscriminate rules allows us to efficiently and soundly prune the hypothesis space and thus improve learning performance. Although existing approaches find reasons why a rule fails on examples, such as having erroneous literals (Muggleton, Lin, and Tamaddoni-Nezhad 2015; Cropper and Hocquette 2024) or rules (Shapiro 1983; Raghothaman et al. 2020), they do not find reducible or indiscriminate rules. We expand on the novelty in Section 2.
@@ 2 Related Work
ILP. Many ILP systems use bottom clauses (Muggleton 1995; Srinivasan 2001) or variants (Ray 2009) to restrict the hypothesis space. The bottom clause of an example is the most specific clause that entails that example. Building a bottom clause can be expensive (Muggleton 1995) and, in the worst case, a learner needs to build a bottom clause for every positive example. Moreover, approaches that use bottom clauses struggle to learn hypotheses with recursion or predicate invention. By contrast, Reducer does not use bottom clauses.
Redundancy in ILP. Fonseca et al. (2004) define self-redundant clauses, similar to our definition of a reducible rule (Definition 8). Their definition does not guarantee that all the refinements (specialisations) of a redundant clause are redundant. By contrast, we prove that specialisations of a reducible rule are reducible (Proposition 1). Moreover, the authors do not propose a way to detect such rules and expect users to provide information about them. By contrast, we introduce Reducer, which automatically finds reducible and indiscriminate rules. Raedt and Ramon (2004) check whether a rule has a redundant atom before testing it on examples to avoid the coverage check. This approach requires anti-monotonic constraints, where if a constraint holds for a rule then it holds for all its generalisations, but not necessarily its specialisations. By contrast, we find properties which allow us to prune specialisations of a rule. Moreover, Raedt and Ramon (2004) do not explicitly identify implications between literals and thus could keep building rules with the same implied literals. By contrast, Reducer explicitly finds implications between literals to prune the hypothesis space. Zeng, Patel, and Page (2014) prune rules with syntactic redundancy. For instance, for the rule h(X) $\leftarrow$ p(X,Y), p(X,Z), the authors detect that p(X,Z) is duplicate to the literal p(X,Y) under the renaming $Z\mapsto Y$, where $Z$ and $Y$ are not in other literals. By contrast, we detect semantic redundancy by finding reducible and indiscriminate rules. Srinivasan and Kothari (2005) reduce the dimensionality of bottom clauses using statistical methods to compress bottom clauses. The resulting lower-dimensional space translates directly to a smaller hypothesis space. We differ by not using bottom clauses.
@@
Rule selection. Many recent systems formulate the ILP problem as a rule selection problem (Corapi, Russo, and Lupu 2011; Kaminski, Eiter, and Inoue 2019; Raghothaman et al. 2020; Bembenek, Greenberg, and Chong 2023; Law 2023). These approaches precompute every possible rule in the hypothesis space and then search for a subset that generalises the examples. Because they precompute all possible rules, they cannot learn rules with many literals. Moreover, because they precompute all possible rules, they can build pointless rules. For instance, if allowed to use the relations int/1 and even/1, ilasp4 (Law 2023) will precompute all possible rules with int(A) and even(A) in the body. By contrast, Reducer does not precompute all rules and instead builds constraints from pointless rules to restrict the generation of rules. For instance, if Reducer sees a rule containing int(A) and even(A), it will identify that even(A) implies int(A) and will henceforth never build a rule with both literals.
@@
Constraints. Many recent ILP systems frame the ILP problem as a constraint satisfaction problem (Corapi, Russo, and Lupu 2011; Ahlgren and Yuen 2013; Schüller and Benz 2018; Kaminski, Eiter, and Inoue 2019; Bembenek, Greenberg, and Chong 2023). MUSPer (Cropper and Hocquette 2024) finds minimal sub-hypotheses that can never be true and builds constraints from them to prune the hypothesis space. For instance, given the rule f(A) $\leftarrow$ succ(A,B),odd(B),even(B), MUSPer can identify that a number cannot be both odd and even and then prunes the hypothesis space accordingly. By contrast, Reducer finds pointless rules with redundancy.
@@
Rule induction. ILP approaches induce rules from data, similar to rule learning methods (Fürnkranz and Kliegr 2015), such as AMIE+ (Galárraga et al. 2015) and RDFRules (Zeman, Kliegr, and Svátek 2021). Most rule-mining methods are limited to unary and binary relations, require facts as input, and operate under an open-world assumption. By contrast, Reducer operates under a closed-world assumption, supports relations of any arity, and can learn from definite programs as background knowledge.
@@
Redundancy in AI. Theorem proving has preprocessing techniques for eliminating redundancies (Hoder et al. 2012; Khasidashvili and Korovin 2016; Vukmirovic, Blanchette, and Heule 2023). In SAT, redundancy elimination techniques, such as blocked clause elimination  (Kullmann 1999), play an integral role in modern solvers  (Heule, Järvisalo, and Biere 2010; Barnett, Cerna, and Biere 2020; Biere, Järvisalo, and Kiesl 2021). Similar to this paper, the goal of redundancy elimination methods in SAT is to identify clauses in a SAT input containing literals that always resolve to tautologies.
@@ 3 Problem Setting
We assume familiarity with logic programming (Lloyd 2012) but have included summaries in the supplementary material. For clarity, we define some key terms. A rule $r$ is a definite clause of the form $h\leftarrow p_{1},\ldots,p_{n}$ where $h,p_{1},\ldots,p_{n}$ are literals, $head(r)=h$, and $body(r)=\{p_{1},\ldots,p_{n}\}$. We denote the set of variables in a literal $l$ as $vars(l)$. The variables of a rule $r$, denoted $vars(r)$, is defined as $vars(head(r))\cup\bigcup_{p\in body(r)}vars(p)$. Given a rule $r$ and a set of positive literals $C$, by $r\cup C$ we denote the rule $r^{\prime}$ such that $head(r)=head(r^{\prime})$ and $body(r^{\prime})=body(r)\cup C$. A specialisation of a rule $r$ is a rule $r^{\prime}$ that is $\theta$-subsumptively more specific than $r$ (See (Plotkin 1971) and the Appendix). We focus on a restricted form of $\theta$-subsumption which only considers whether the bodies of two rules with the same head literal are contained in one another. We formally define this restriction through a subrule relation:
@@
The sub-hypothesis relation captures a particular type of hypothesis we refer to as basic. These are hypotheses for which specific rules do not occur as part of a recursive predicate definition:
@@
As we show in Section 3, certain rules allow us to prune hypotheses basic with respect to these rules.
@@
Given an ILP input $(E,B,\mathcal{H})$, a cost function $cost_{E,B}~:$$\mathcal{H}~\mapsto$$\mathbb{N}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
@@
We want to find pointless rules, i.e., rules that cannot be in an optimal hypothesis. We focus on reducible and indiscriminate rules.
@@
The rule r1 is reducible because odd(A) implies int(A). Therefore, r1 is logically equivalent r2.
@@
$r_{2}$ = h $\leftarrow$ member(L,X), member(L,Y), X>>Y
@@
Let $r_{1}$ be a rule, $r_{2}\subseteq r_{1}$, $l\in body(r_{2})$, and $l$ be $r_{2}$-captured. Then $l$ is $r_{1}$-captured.
@@
Let $r$ be a rule, $B$ be BK, $l\in body(r)$ be $r$-captured, and $B\models r\leftrightarrow r\setminus\{l\}$. Then $r$ is reducible.
@@
Let $l$ be a $r_{1}$-captured literal and $B\models r_{1}\leftrightarrow(r_{1}\setminus\{l\})$. By Lemma 1, $l$ is also $r_{2}$-captured. Let $C=body(r_{2})\setminus body(r_{1})$. Then $B\models(r_{1}\cup C)\leftrightarrow((r_{1}\cup C)\setminus\{l\}).$ Finally, since $r_{2}=(r_{1}\cup C)$, we deduce $B\models r_{2}\leftrightarrow(r_{2}\setminus\{l\})$. ∎
@@
By Proposition 1, $r_{1}$ is also reducible implying that there exists a rule $r_{3}\subseteq r_{1}$ such that (i) $B\models r_{1}\leftrightarrow r_{3}$ and (ii) $|r_{3}|<|r_{1}|$. Let $h_{3}=(h_{1}\setminus\{r_{1}\})\cup\{r_{3}\}$. Then $cost_{E,B}(h_{3})<cost_{E,B}(h_{1})$, i.e. $h_{1}$ is not optimal. ∎
@@
We introduce indiscriminate rules. To motivate them, consider the rule:
@@
This rule is not reducible because odd(A) does not imply lt(A,10), nor does lt(A,10) imply odd(A). However, suppose we have the negative examples $E^{-}=\{f(1),f(2),f(3)\}$. For these examples, the literal lt(A,10) implies all the negative examples. In other words, there is no negative example greater than 10. Therefore, this literal, and thus this rule, is pointless because it cannot discriminate against the negative examples. We formalise this notion of an indiscriminate rule:
@@
Let $r$ be a rule, $B$ be BK, $E^{-}$ be negative examples with the same predicate symbol as the head of $r$, $l\in body(r)$ be $r$-captured, and for all $e\in E^{-}$, $B\models(r\rightarrow e)\leftrightarrow(r\setminus\{l\}\rightarrow e)$. Then $r$ is indiscriminate.
@@
Let $l$ be an $r_{1}$-captured literal and for all $e\in E^{-}$, $B\models(r_{1}\rightarrow e)\leftrightarrow(r_{1}\setminus\{l\}\rightarrow e)$. By Lemma 1, $l$ is also $r_{2}$-captured. Let $C=body(r_{2})\setminus body(r_{1})$. We deduce the following for all $e\in E^{-}$:
$B\models eq(r_{1},e)\rightarrow eq(r_{1}\cup C,e)$
where $eq(r^{\prime},e)\equiv(r^{\prime}\rightarrow e)\leftrightarrow(r^{\prime}\setminus\{l\}\rightarrow e).$ Since $r_{2}=(r_{1}\cup C)$ then for all $e\in E^{-}$, $B\models(r_{2}\rightarrow e)\leftrightarrow(r_{2}\setminus\{l\}\rightarrow e)$. ∎
@@
By Proposition 3, $r_{1}$ is also reducible implying that there exists a rule $r_{3}\subseteq r_{1}$ such that (i) for all $e\in E^{-}$, $B\models(r_{1}\rightarrow e)\leftrightarrow(r_{3}\rightarrow e)$ and (ii) $|r_{3}|<|r_{1}|$. Let $h_{3}=(h_{1}\setminus\{r_{1}\})\cup\{r_{3}\}$. Then $cost_{E,B}(h_{3})<cost_{E,B}(h_{1})$, i.e. $h_{1}$ is not optimal. ∎
@@ 4 Algorithm
We now describe Reducer (Algorithm 1), which builds on Popper by identifying pointless rules and using them to prune the hypothesis space. Reducer takes as input background knowledge (bk), positive (pos) and negative (neg) training examples, and a maximum hypothesis size (max_size). Reducer follows Popper and uses a generate, test, and constrain loop to find an optimal hypothesis (Definition 6). Reducer starts with an answer set program $\mathcal{P}$. The (stable) models of $\mathcal{P}$ correspond to hypotheses (definite programs) and represent the hypothesis space. In the generate stage (line 6), Reducer uses an answer set programming (ASP) solver to find a model of $\mathcal{P}$. If there is no model, Reducer increments the hypothesis size and loops again (lines 7-9). If there is a model, Reducer converts it to a hypothesis $h$. In the test stage, it uses Prolog to test $h$ on the training examples (line 10). If $h$ is better (according to a cost function) than the previous best-seen hypothesis, Reducer sets the best hypothesis to $h$ (line 12). In the constrain stage, Reducer builds hypothesis constraints (represented as ASP constraints) from $h$. This step is identical to Popper. For instance, if $h$ does not entail any positive examples, Reducer builds a specialisation constraint to prune its specialisations. Reducer saves these constraints (line 13) which it passes to the generate stage to add to $\mathcal{P}$ to prune models and thus prune the hypothesis space.
@@
The novelty of Reducer is checking whether a hypothesis has a pointless rule (line 14) by calling Algorithm 2. We describe Algorithm 2 below. If a hypothesis has a pointless rule, Reducer builds constraints to prune its generalisations and specialisations, in which the pointless rule is basic.
@@
Algorithm 2 checks whether a hypothesis has a rule which (i) is basic111 We enforce other syntactic restrictions, such as forcing literals in a rule to be connected, where they cannot be partitioned into two sets such that the variables in the literals of one set are disjoint from the variables in the literals of the other set. (Definition 3), (ii) has a captured literal (Definition 7, line 8), and (iii) is reducible (Definition 8, line 10) or indiscriminate (Definition 9, line 12). Algorithm 2 includes two subprocedures for checking if a rule is reducible or indiscriminate. The first procedure (lines 16-18) checks if the query $\mbox{body}^{\prime}\cup\{\neg\mbox{literal}\}$ is unsatisfiable over the background knowledge. This query determines reducibility because it checks whether a literal is implied by the body. The second subprocedure (lines 20-24) checks if the subrule without the captured literal entails the same negative examples as the original rule. This query matches Definition 9. We use Prolog to perform the unsat (line 18) and coverage (lines 22-23) checks.
@@
Cropper and Morel (2021) show that given optimally sound constraints Popper returns an optimal hypothesis if one exists (Theorem 1). Reducer builds on Popper by pruning pointless (one with a reducible or indiscriminate rule) hypotheses. By Corollary 1, a pointless hypothesis is not optimal. Therefore, Reducer never prunes an optimal hypothesis so returns one if it exists. ∎

### v3

@@ Abstract
Codehttps://github.com/logicand-learning-lab/aaai26-implications
@@ 1 Introduction
Given this input, we might want to learn a rule such as:
@@
However, $r_{2}$ and $r_{3}$ are pointless because odd(A) implies int(A), so we can remove int(A) whilst preserving semantics. We call such rules reducible rules.
@@
The novelty of this paper is to show that ignoring reducible and indiscriminate rules allows us to efficiently and soundly prune the hypothesis space and thus improve learning performance. Although existing approaches find reasons why a rule fails on examples, such as having erroneous literals (Muggleton et al. 2015; Cropper and Hocquette 2024) or rules (Shapiro 1983; Raghothaman et al. 2020), they do not find reducible or indiscriminate rules.
@@ 2 Related Work
ILP. Many ILP systems use bottom clauses (Muggleton 1995; Srinivasan 2001) or variants (Ray 2009) to restrict the hypothesis space. Building bottom clauses can be expensive (Muggleton 1995). In the worst case, a learner needs a bottom clause for every positive example. Bottom clause approaches struggle to learn hypotheses with recursion or predicate invention. Reducer does not use bottom clauses.
Redundancy in ILP. Fonseca et al. (2004) define self-redundant clauses similar to reducible rules (Definition 8) but cannot guarantee that specialisations are redundant (unlike Proposition 1) and require users to provide redundancy information. Raedt and Ramon (2004) check redundancy before testing but require anti-monotonic constraints (holding for generalisations, not specialisations) and do not identify implications between literals. By contrast, Reducer finds implications to prune specialisations. Zeng et al. (2014) prune syntactic redundancy (e.g. duplicate variables), whereas we detect semantic redundancy. Srinivasan and Kothari (2005) compress bottom clauses statistically; we avoid bottom clauses.
@@
Rule selection. Many systems (Corapi et al. 2011; Kaminski et al. 2019; Raghothaman et al. 2020; Bembenek et al. 2023; Law 2023) precompute every possible rule in the hypothesis space and then search for a subset that generalises the examples. Because they precompute all possible rules, they cannot learn rules with many literals and can build pointless rules. For instance, ilasp4 (Law 2023) will precompute all rules with int(A) and even(A) in the body. By contrast, Reducer builds constraints from pointless rules to restrict rule generation. If Reducer sees a rule containing int(A) and even(A), it identifies that even(A) implies int(A) and henceforth never builds a rule with both literals.
@@
Constraints. Many recent ILP systems frame the ILP problem as a constraint satisfaction problem (Corapi et al. 2011; Ahlgren and Yuen 2013; Schüller and Benz 2018; Kaminski et al. 2019; Bembenek et al. 2023). MUSPer (Cropper and Hocquette 2024) finds minimal sub-hypotheses that can never be true and builds constraints from them to prune the hypothesis space. For instance, given the rule f(A) $\leftarrow$ succ(A,B),odd(B),even(B), MUSPer can identify that a number cannot be both odd and even and then prunes the hypothesis space accordingly. By contrast, Reducer finds pointless rules with redundancy.
@@
Rule induction. ILP approaches induce rules from data, similar to rule learning methods (Fürnkranz and Kliegr 2015) such as AMIE+ (Galárraga et al. 2015) and RDFRules (Zeman et al. 2021). Most rule-mining methods are limited to unary and binary relations, require facts as input, and operate under an open-world assumption. By contrast, Reducer operates under a closed-world assumption, supports arbitrary-arity relations, and can learn from definite programs as background knowledge.
@@
Redundancy in AI. Theorem proving has preprocessing techniques for eliminating redundancies (Hoder et al. 2012; Khasidashvili and Korovin 2016; Vukmirovic et al. 2023). In SAT, redundancy elimination techniques, such as blocked clause elimination (Kullmann 1999), play an integral role in modern solvers (Heule et al. 2010; Barnett et al. 2020; Biere et al. 2021). Similar to this paper, SAT redundancy elimination identifies clauses containing literals that always resolve to tautologies.
@@ 3 Problem Setting
We assume familiarity with logic programming (Lloyd 2012) but have included summaries in the supplementary material. For clarity, we define some key terms. A rule $r$ is a definite clause of the form $h\leftarrow p_{1},\ldots,p_{n}$ where $h,p_{1},\ldots,p_{n}$ are literals, $head(r)=h$, and $body(r)=\{p_{1},\ldots,p_{n}\}$. A definite program is a set of definite clauses. We use the term hypothesis interchangeably with a definite program. We denote the set of variables in a literal $l$ as $vars(l)$. The variables of a rule $r$, denoted $vars(r)$, is defined as $vars(head(r))\cup\bigcup_{p\in body(r)}vars(p)$. Given a rule $r$ and a set of literals $C$, by $r\cup C$ we denote the rule $r^{\prime}$ such that $head(r^{\prime})=head(r)$ and $body(r^{\prime})=body(r)\cup C$. A specialisation of a rule $r$ is a rule $r^{\prime}$ that is $\theta$-subsumptively more specific than $r$, where $\theta=\{v_{1}/t_{1},…,v_{n}/t_{n}\}$ is a substitution that simultaneously replaces each variable $v_{i}$ by term $t_{i}$ (See (Plotkin 1971) and the Appendix). We focus on a restricted form of $\theta$-subsumption which only considers whether the bodies of two rules with the same head literal are contained in one another, i.e., no variable renaming. We formally define this restriction through a subrule relation:
@@
The sub-hypothesis relation captures a particular type of hypothesis that we refer to as basic. These are hypotheses for which specific rules do not occur as part of a recursive predicate definition:
@@
As we show in Section 3, under certain conditions we can prune hypotheses containing sub-hypotheses that are basic with respect to a contained rule.
@@
Given an ILP input $(E,B,\mathcal{H})$, a cost function $cost_{E,B}~:$$\mathcal{H}~\to$$\mathbb{N}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
@@
We want to find rules that cannot be in an optimal hypothesis. We focus on reducible and indiscriminate rules.
@@
The rule r1 is reducible because odd(A) implies int(A). Therefore, r1 is logically equivalent to r2.
@@
$r_{2}$ = h $\leftarrow$ member(L,X), member(L,Y), gt(X,Y)
@@
Let $r_{1}$ and $r_{2}$ be rules such that $r_{2}\subseteq r_{1}$, $l\in body(r_{2})$, and $l$ be $r_{2}$-captured. Then $l$ is $r_{1}$-captured.
@@
Let $r$ be a rule, $B$ be BK, $l\in body(r)$ be $r$-captured, and $B\models(body(r)\setminus\{l\})\rightarrow l$. Then $r$ is reducible.
@@
Let $l$ be a $r_{1}$-captured literal and $B\models(body(r_{1})\setminus\{l\})\rightarrow l$. By Lemma 1, $l$ is also $r_{2}$-captured. Let $C=body(r_{2})\setminus body(r_{1})$. Then, $(body(r_{1})\setminus\{l\})\rightarrow l$ is subsumptively more general than $(body(r_{1}\cup C)\setminus\{l\})\rightarrow l$, where $r_{2}=r_{1}\cup C$, i.e., if the former holds for $r_{1}$ than the latter holds for $r_{2}$. Thus, $r_{2}$ is reducible. ∎
@@
By Proposition 1, $r_{1}$ is also reducible implying that there exists an $l\in body(r_{1})$ and rule $r_{3}\subset r_{1}$ such that (i) $B\models(body(r_{1})\setminus\{l\})\rightarrow l$, (ii) $r_{1}=r_{3}\cup\{l\}$, and (iii) $|r_{3}|<|r_{1}|$. Furthermore, $r_{3}$ is basic given that $r_{1}$ is basic. Let $h_{3}=(h_{1}\setminus\{r_{1}\})\cup\{r_{3}\}$. Then $cost_{E,B}(h_{3})<cost_{E,B}(h_{1})$, i.e. $h_{1}$ is not optimal. ∎
@@
We introduce indiscriminate rules, a weakening of reducible rules. To motivate them, consider the rule:
@@
This rule is not reducible because odd(A) does not imply lt(A,10), nor does lt(A,10) imply odd(A). However, suppose we have the negative examples $E^{-}=\{f(1),f(2),f(3)\}$. For these examples, the literal lt(A,10) implies all the negative examples. In other words, there is no negative example greater than 10. Therefore, this literal (and thus this rule) is pointless because it cannot discriminate against the negative examples. We formalise this notion of an indiscriminate rule:
@@
Let $r$ be a rule, $B$ be BK, $E^{-}$ be negative examples with the same predicate symbol as the head of $r$, $l\in body(r)$ be $r$-captured, and for all $e\in E^{-}$, $B\models(body(r)\setminus\{l\})\theta_{e}\rightarrow l\theta_{e}^{r}$ where $\theta_{e}^{r}$ is a substitution with domain $vars(head(r))$ such that $head(r)\theta_{e}^{r}=e$. Then $r$ is indiscriminate.
@@
Let $l$ be an $r_{1}$-captured literal and for all $e\in E^{-}$, $B\models(body(r_{1})\setminus\{l\})\theta_{e}^{r_{1}}\rightarrow l\theta_{e}^{r_{1}}$. By Lemma 1, $l$ is also $r_{2}$-captured. Let $C=body(r_{2})\setminus body(r_{1})$. We deduce the following for all $e\in E^{-}$: $(body(r_{1})\setminus\{l\})\theta_{e}^{r_{1}}\rightarrow l\theta_{e}^{r_{1}}$ is subsumptively more general than $(body(r_{1}\cup C)\setminus\{l\})\theta_{e}^{r_{2}}\rightarrow l\theta_{e}^{r_{2}}$ where $r_{2}=(r_{1}\cup C)$, i.e., if the former holds for $r_{1}$, than the latter holds for $r_{2}$. Thus, $r_{2}$ is indiscriminate. ∎
@@
By Proposition 3, $r_{1}$ is also indiscriminate implying that there exists $l\in body(r_{1})$ and $r_{3}\subset r_{1}$ such that (i) for all $e\in E^{-}$, $B\models(body(r_{1})\setminus\{l\})\theta_{e}^{r_{1}}\rightarrow l\theta_{e}^{r_{1}}$, (ii) $r_{1}=r_{3}\cup\{l\}$, and (iii) $|r_{3}|<|r_{1}|$. Furthermore, $r_{3}$ is basic given that $r_{1}$ is basic. Let $h_{3}=(h_{1}\setminus\{r_{1}\})\cup\{r_{3}\}$. Then $cost_{E,B}(h_{3})<cost_{E,B}(h_{1})$, i.e. $h_{1}$ is not optimal. ∎
@@ 4 Algorithm
We now describe Reducer (Algorithm 1), which builds on Popper by identifying pointless rules and using them to prune the hypothesis space. Reducer takes as input background knowledge (bk), positive (pos) and negative (neg) training examples, and a maximum hypothesis size (max_size). Reducer follows Popper and uses a generate, test, and constrain loop to find an optimal hypothesis (Definition 6). Reducer starts with an answer set program $\mathcal{P}$. The (stable) models of $\mathcal{P}$ correspond to hypotheses (definite programs) and represent the hypothesis space. In other words, $\mathcal{P}$ encodes in ASP the search for a syntactically valid hypothesis that satisfies given constraints. In the generate stage (line 6), Reducer uses an answer set programming (ASP) solver to find a model of $\mathcal{P}$. If there is no model, Reducer increments the hypothesis size and loops again (lines 7-9). If there is a model, Reducer converts it to a hypothesis $h$. In the test stage, it uses Prolog to test $h$ on the training examples (line 10). If $h$ is better (according to a cost function) than the previous best-seen hypothesis, Reducer sets the best hypothesis to $h$ (line 12). In the constrain stage, Reducer builds hypothesis constraints (represented as ASP constraints) from $h$. This step is identical to Popper. For instance, if $h$ does not entail any positive examples, Reducer builds a specialisation constraint to prune its specialisations. Reducer saves these constraints (line 13) which it passes to the generate stage to add to $\mathcal{P}$ to prune models and thus prune the hypothesis space.
@@
The novelty of Reducer is checking whether a hypothesis has a pointless rule (line 14) by calling Algorithm 2. We describe Algorithm 2 below. If a hypothesis has a pointless rule, Reducer builds constraints to prune its generalisations and specialisations, in which the pointless rule is basic. Reducer passes these constraints to the generate stage to add to $\mathcal{P}$ to prune models and thus prune the hypothesis space of other hypotheses with pointless rules.
@@
Algorithm 2 checks whether a hypothesis has a rule which (i) is basic [^1] We enforce other syntactic restrictions, such as forcing literals in a rule to be connected, where they cannot be partitioned into two sets such that the variables in the literals of one set are disjoint from the variables in the literals of the other set. (Definition 3), (ii) has a captured literal (Definition 7, line 8), and (iii) is reducible (Definition 8, line 10) or indiscriminate (Definition 9, line 12). Algorithm 2 includes two subprocedures for checking if a rule is reducible or indiscriminate. The first procedure (lines 16-18) checks if the query $\mbox{body}^{\prime}\cup\{\neg\mbox{literal}\}$ is unsatisfiable over the background knowledge. This query determines reducibility because it checks whether a literal is implied by the body. The second subprocedure (lines 20-24) checks if the subrule without the captured literal entails the same negative examples as the original rule. This query matches Definition 9. We use Prolog to perform the unsat (line 18) and coverage (lines 22-23) checks.
@@
Cropper and Morel (2021) show that given optimally sound constraints (an optimally sound constraint never prunes an optimal hypothesis), Popper returns an optimal hypothesis if one exists (Theorem 1). Reducer builds on Popper by pruning pointless (one with a reducible or indiscriminate rule) hypotheses. By Corollary 1, a pointless hypothesis is not optimal. Therefore, Reducer never prunes an optimal hypothesis so returns one if it exists. ∎